OpenStudy (anonymous):

y = -9x^2 y^2 = 3x find y and x

OpenStudy (anonymous):

you still here?

jhonyy9 (jhonyy9):

y= -9x2 y2= 3x so from this y= +/- sqrt(3x) than we get that -9x2= +/- sqrt(3x)t what if squared result 81x4= 3x divide by 3x and we get 27x3 =1 x3= 1/27 x3= 3 on exponent (-3) x= (3 on exponent (-3)) on exponent 1/3 x= 3 on exponent (-3*1/3) = 3 on exponent (-1) x=1/3 so than y= -9(1/9) = -9/9 = -1 x= 1/3 y= -1

jhonyy9 (jhonyy9):

this is understandably ?

OpenStudy (anonymous):

i got it =D thank you

jhonyy9 (jhonyy9):

1/27 = 1/3 on exponent 3 because 3 on exponent 3 is 27 so but because 27 is on denominator so for there is 3 on exponent minus 3 = 1/27

jhonyy9 (jhonyy9):

right ?

jhonyy9 (jhonyy9):

so when we h

jhonyy9 (jhonyy9):

we have one term cubed equal a so like x on exponent 3 = a than x= cube root x what equal x on exponent 1/3

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

and if we have one term x on exponent a and this all on exponent b so than we get xon exponent a*b

jhonyy9 (jhonyy9):

x on exponent a * b

jhonyy9 (jhonyy9):

ok ???

OpenStudy (anonymous):

yeahh i got it now :) i understand you steps thank you so much <3

jhonyy9 (jhonyy9):

yw